OpenStudy (anonymous):

solve by elimination; -8x - 20y = -4 6 x + 10y = -12 ---------------------------- 24y = 12 x - 24 8x = 16y + 16

OpenStudy (amistre64):

2* bottom will eliminate the ys

OpenStudy (amistre64):

the other one should be put into a format where they at least match somewaht to see it better

OpenStudy (amistre64):

divide it all by 4 to lessen the numbers 6y = 3x - 6 2x = 4y + 4 you sure that aint typoed?

OpenStudy (anonymous):

-8x - 20y = -4 6x + 10y = -12

OpenStudy (anonymous):

or..

OpenStudy (anonymous):

24y = 12x - 24 8x = 16y + 16